# Predictive Linear Online Tracking for Unknown Targets

### v1

@@ Appendix F Implementation on Quadrotors
The online tracking setting is of particular interest when applied to the control of quadrotors to track either a virtual target or another object of interest. We extend the simulation setup of Appendix E to a practical one by applying PLOT to track a virtual sequentially revealed target using a Crazyflie 2.1 Quadrotor Bitcraze [2023b] depicted in Figure 15. The controllers use the linearised model of the quadrotor and the derived in Appendix E.1. Below, we detail our experimental setup, as well as the results of the online tracking experiments.
@@
Figure 14: The Crazyflie 2.1 Quadrotor with three visual markers.
Figure 15: The Setup for Flying Experiments. There is a dedicated Flying Arena, and $3$ of the $5$ mounted VICON Cameras to locate the quadrotors.
F.1 Experimental Setup
@@
We set up the control architecture with the Robotic Operating System (ROS) Quigley et al. [2009], which allows the controller to receive the state of the drone from the VICON in real-time. The control action is computed on a local computer and communicated with the drone by sending variable-sized packets through a $2.4$Ghz USB dongle Bitcraze [2023a].
The virtual reference trajectory is generated online after the drone has successfully taken off and is at a pre-defined random initial hovering position at a height of $0.4$ meters. The controller receives the virtual target state through the ROS network as it flies at a predetermined rate of $10$ Hz.
@@
Figure 16: Circular reference tracked by Crazyflie Drones with the PLOT and LQR controllers.
@@
Figure 17: “Infinity”-shape reference tracked by Crazyflie Drones with the PLOT and LQR controllers.
F.2 Online Tracking Experiments
We run PLOT on the described setup to show its tracking performance. The cost matrices are taken to be the same as in Appendix E.1, the prediction horizon is fixed to $W=5$ and the forgetting factor to $\gamma=0.8$. The initialization of PLOT is the same as in Appendix E.3.1. Two virtual reference target shapes, a circle and an “infinity”-shape are generated online and published to the ROS network. The quadrotors start in the flying arena in Figure 15 and follow the target as soon as state information is received. Figures 16 and 17 below show the trajectory and state plots of PLOT and Naive LQR implemented for a horizon of $T=40$ seconds for the circle and “infinity”-shape reference trajectories, respectively. Naive LQR is implemented as described in Appendix E.3.1.
As in the simulations, without any affine term, the naive LQR controller exhibits a delayed tracking behavior as expected. The constant offset can be noted clearly for both shapes. In comparison, the PLOT algorithm experiences a smaller tracking error. As opposed to the linear simulations in Appendix E, the nonlinear hardware implementation has additional errors, especially for the circular example, due to imperfect mass value or linearization error, among other reasons. The VICON system is also known to introduce a drift through its localization algorithm when some of the cameras are malfunctioning.

### v3

@@ Appendix G Implementation on Quadrotors
The online tracking setting is of particular interest when applied to the control of quadrotors to track either a virtual target or another object of interest. We extend the simulation setup of Appendix F to a practical one by applying PLOT to track a virtual sequentially revealed target using a Crazyflie 2.1 Quadrotor (Bitcraze, 2023b) depicted in Figure 21. The controllers use the linearised model of the quadrotor and the derived in Appendix F.1. Below, we detail our experimental setup, as well as the results of the online tracking experiments.
@@
Figure 21: The D-FALL Laboratory of ETH Zürich, consisting of a dedicated Flying Arena and $5$ ($3$ shown) mounted VICON Cameras for state estimation.
G.1 Experimental Setup
@@
We set up the control architecture with the Robotic Operating System (ROS) (Quigley et al., 2009), which allows the controller to receive the state of the drone from the VICON mocap system in real-time. The control action is computed on a local computer and communicated with the drone by sending variable-sized packets through a $2.4$Ghz USB dongle (Bitcraze, 2023a). The virtual reference trajectory is generated online after the drone has successfully taken off and is at a pre-defined random initial hovering position at a height of $0.4$ meters. The controller receives the target state through the ROS network as it flies at a predetermined rate of $10$ Hz.
G.2 Online Tracking Experiments
We run PLOT on the described setup to show its tracking performance. The cost matrices are taken to be the same as in Appendix F.1, the prediction horizon is fixed to $W=5$ for the “infinity”-shape and $W=3$ for the circle, the forgetting factor to $\gamma=0.8$ and no projection is performed. The initialization of PLOT is the same as in Appendix F.3.1. Two virtual reference target shapes, a circle and an “infinity”-shape are generated online and published to the ROS network. The quadrotors start in the flying arena in Figure 21 and follow the target as soon as state information is received. Figures 22 and 23 below show the trajectory and state plots of PLOT and Naive LQR implemented for a horizon of $T=40$ seconds for the circle and “infinity”-shape reference trajectories, respectively. Naive LQR is implemented as described in Appendix F.3.1.
As in the simulations, without any affine term, the naive LQR controller exhibits a delayed tracking behavior as expected. The constant offset can be noted clearly for both shapes. In comparison, the PLOT algorithm experiences a smaller tracking error. As opposed to the linear simulations in Appendix F, the nonlinear hardware implementation has additional errors, especially for the circular example, due to imperfect mass value or linearization error, among other reasons. The VICON system is also known to introduce a drift through its localization algorithm when some of the cameras are malfunctioning.
@@ G.2 Online Tracking Experiments
Figure 22: Circular reference tracked by Crazyflie Drones with the PLOT and Naive LQR controllers.
@@
Figure 23: “Infinity”-shaped reference tracked by Crazyflie Drones with the PLOT and Naive LQR controllers.